find all numbers "x" that satisfy the given condition: x^2 + 10x + 16 = e^3
Answer by edjones(8007)

x^2 + 10x + 16 = e^3
x^2+10x =e^3-16
x^2+10x+25=e^3+9 completing the square.
(x+5)^2=e^3+9
x+5=+-sqrt(e^3+9)
x=-5+sqrt(e^3+9), x=-5-sqrt(e^3+9)
.
Ed